Solve each system of equations of substitution.
3x-7y=-6
x+2y=11

Answer by ReadingBoosters(3246)

3x-7y=-6
x+2y=11 or x = 11-2y
Substitute
3(11-2y) - 7y = -6
Distribute the 3
33 - 6y - 7y = -6
-13y + 33 = -6
-13y = -39
-13y/-13 = -39/-13
y = 3
x = 11 - 2(3) = 5